OpenStudy (anonymous):

The relationship between the sale price of an item, S, the list price L, and the discount rate r is related by the formula S = L(1 – r). Your friend Albert bought a coat on sale for $72 that had a list price of $120. Solve for r and find the discount rate. Answer r = 1 – S/L, 40% r = 1 + S/L, 60% r = 1 – S/L, 60% r = 1 + S/L, 40%

OpenStudy (anonymous):

We have the equation \[S=L(1-r)\] Divide both sides by L, and get: \[S/L = 1-r\] subtract 1 from both sides: \[S/L - 1 = -r\] Multiply both sides by -1: \[1- S/L = r\] Now plug in the values from the question to solve for r.

OpenStudy (anonymous):

its either a or c

OpenStudy (anonymous):

The list price is $120 and the sale price is $72. Simply plug into the equation.